OpenStudy (anonymous):

Showing all of your work, solve for x: 4x - 6 = x + 9

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

plug it in

OpenStudy (welshfella):

first subtract x from both sides of the equation then add 6 to both sides then divide both sides by 3

OpenStudy (welshfella):

@marinesrule - please don't just give answers. Read Code of Conduct

OpenStudy (anonymous):

So would I just put x=5??

OpenStudy (anonymous):

sorry

OpenStudy (welshfella):

No you need to show your workk I'll do the first line for you 4x - x - 6 = x - x + 9 3x - 6 = 9

OpenStudy (anonymous):

i need help wats 1-1

OpenStudy (welshfella):

note that by subtracting x from both sides the x on the right disappears so to get rid of the -6 on the left we add 6 to both sides 3x - 6 + 6 = 9 + 6 3x = 15

OpenStudy (welshfella):

now to get the value of x we divide both sides by 3